OpenStudy (some.random.cool.kid):

The graph of y = 3x2 + 4x − 7 is shown below. What are the zeros of the function (as exact values), the y-intercept, and the maximum or minimum value of the function (rounded to the nearest tenth)? Hint: Use a graphing utility or rewrite the function in different forms to help you answer these questions. The graph of the parabolic function y equals three times x squared plus four times x minus seven

OpenStudy (some.random.cool.kid):

Zeros: x equals negative start fraction seven over three end fraction comma 1; y-intercept: (0, −7); minimum: −8.3 Zeros: x equals start fraction three over seven end fraction comma −1; y-intercept: (0, −7); maximum: −8.3 Zeros: x equals start fraction seven over three end fraction comma −1; y-intercept: (0, 7); maximum: −0.7 Zeros: x equals negative start fraction three over seven end fraction comma 1; y-intercept: (0, 7); minimum: −0.7

OpenStudy (some.random.cool.kid):

@Nnesha

OpenStudy (some.random.cool.kid):

@ParthKohli @sammixboo @sleepyhead314

OpenStudy (some.random.cool.kid):

@Data_LG2 @freckles @timo86m

OpenStudy (maddy1251):

I can help.

OpenStudy (some.random.cool.kid):

@Maddy1251 think you can help some more I was able to take a break after finishing two assignments so now I'm back with a different one.

OpenStudy (some.random.cool.kid):

oh your backlol

OpenStudy (some.random.cool.kid):

welcome to 60 ss btw

OpenStudy (maddy1251):

Thank ya :-)

OpenStudy (maddy1251):

I pointed out your zeros.

OpenStudy (maddy1251):

x=-2.333 and x=1

OpenStudy (some.random.cool.kid):

ok

OpenStudy (maddy1251):

x=-7/3 instead of -2.333

OpenStudy (maddy1251):

A is your answer.

OpenStudy (maddy1251):

y=3x^2+4x-7 That -7 at the end tells us what our y-intercept is, y=(0,-7)

OpenStudy (some.random.cool.kid):

:o that was fast and ok yeah I was just reading what you wrote.

OpenStudy (maddy1251):

It has a minumum point of x=-0.667 y=-8.333

OpenStudy (maddy1251):

Zeros: x equals negative start fraction seven over three end fraction comma 1; y-intercept: (0, −7); minimum: −8.3 Looks like this matches pretty good.

OpenStudy (some.random.cool.kid):

ahhh thanks I was about to ask how to get 8.3 but you just explained it so... !D

OpenStudy (maddy1251):

Yes, its a parabola, so the bottom of the parabola is our minimums. You'll learn more of that in calc.

OpenStudy (maddy1251):

It would have maximums if the parabola was flipped, I believe, such as opening downwards.

OpenStudy (maddy1251):

Where the U of the parabola would be in the positive quadrants. I could be wrong, I'm dusting the cobwebs off my algebra skills folder today, haha.

OpenStudy (some.random.cool.kid):

haha ok ...

OpenStudy (maddy1251):

Hope that all makes sense.

OpenStudy (some.random.cool.kid):

Ill be back on later I dont want to hold you up ... i have to do the dishes real quick. But i read your stuff and i think i get it.

OpenStudy (maddy1251):

Great, you're not holding me up :-) Just want to make sure you get it. If not let me know

OpenStudy (some.random.cool.kid):

@Maddy1251 yeah thats right im back! the kid returns lol... to much... yeah to much i need a better take away on my intro xD. anyways ready to help me do a few more. After these next three your free and can have a happy day. :`>

OpenStudy (maddy1251):

No your intro is better than anything I could make up xD Yes of course I will help, I love to help you!! Please, as many questions as you need hep with I am here to help. :-)

OpenStudy (some.random.cool.kid):

well that sounds fantastic because I actually had four more but i was going to be nice and do the last one myself.

OpenStudy (some.random.cool.kid):

What is the equation of the quadratic function graphed below? Graph of a parabolic function coming up from the left through the point approximately at negative two and three fourths comma zero and turning at the point negative one comma nine and going down through the point zero comma six and through the point at approximately three fourths comma zero and continuing towards negative infinity. y = 3(x − 1)2 − 9 y = −3(x + 1)2 + 9 y = −3(x − 1)2 − 9 y = 3(x − 9)2 − 1

OpenStudy (some.random.cool.kid):

so this one is pretty similar too just different solutions.

OpenStudy (some.random.cool.kid):

theres my graph.

OpenStudy (maddy1251):

Thank you for the picture :D

OpenStudy (maddy1251):

Its opening downward, so our slope will be Negative. So that takes out A and D

OpenStudy (maddy1251):

Our y-intercept looks to be 6 on the positive side of the y-axis.

OpenStudy (some.random.cool.kid):

yes just about to say that.

OpenStudy (maddy1251):

Soo I am going to simplify out B and C to see if any of these two equations will give me my desired answer.

OpenStudy (maddy1251):

Excellent!

OpenStudy (maddy1251):

I shall start with this one ; y = −3(x + 1)2 + 9

OpenStudy (some.random.cool.kid):

hmm so can I just plug in each equation? or something else?

OpenStudy (some.random.cool.kid):

oh ok

OpenStudy (maddy1251):

I want it to look like ; y=mx+b

OpenStudy (some.random.cool.kid):

3x + 1^2 + 9??

OpenStudy (some.random.cool.kid):

well actually 3x + 3^2 +9* not 1 because I think I distribute both.

OpenStudy (maddy1251):

I have one question

OpenStudy (maddy1251):

y = −3(x − 1)2 − 9 Does it actually look like this? \[y=-3(x-1)^{2} -9\]

OpenStudy (some.random.cool.kid):

yes

OpenStudy (maddy1251):

Ah okay, no wonder why I wasn't getting any right answers!

OpenStudy (some.random.cool.kid):

:O

OpenStudy (maddy1251):

y = −3(x − 1)2 − 9 -3(x-1)(x-1)-9

OpenStudy (some.random.cool.kid):

yep yep

OpenStudy (maddy1251):

(x-1)(x-1)=\[x^{2}-1x-1x+1\]

OpenStudy (some.random.cool.kid):

like terms? do anything with them?

OpenStudy (maddy1251):

\[x^{2}-2x+1\]

OpenStudy (maddy1251):

Look right?

OpenStudy (some.random.cool.kid):

ah you did do something with the like terms. ok proceed, and yeah so far.

OpenStudy (maddy1251):

:-)\[y=-3(x^{2}-2x+1)-9\]

OpenStudy (maddy1251):

Look okay?

OpenStudy (some.random.cool.kid):

yep, yep.

OpenStudy (maddy1251):

\[y=-3x^{2}-6x-3-9\]

OpenStudy (some.random.cool.kid):

oh alright.

OpenStudy (maddy1251):

One minute.. As of right now its in perfect parabola form.Combine like terms for me please.

OpenStudy (some.random.cool.kid):

since there all negatives this will mostly be positives i guess.

OpenStudy (some.random.cool.kid):

I think I figured it out but please continue.

OpenStudy (some.random.cool.kid):

oh okay didnt see that part well just give me a minute.

OpenStudy (some.random.cool.kid):

ok so the 3 and 9 are like terms so -3 -9 = -12?

OpenStudy (some.random.cool.kid):

so it would like this now? 3x^2 - 6x - 12? yes or did I mess up.

OpenStudy (some.random.cool.kid):

I assume i messed up because its suppose to end with a 9. negative or positive but still a 9 and I got a 12.

OpenStudy (some.random.cool.kid):

you still there @Maddy1251

OpenStudy (maddy1251):

Just got back

OpenStudy (maddy1251):

Yes it will be -12

OpenStudy (some.random.cool.kid):

oh ok

OpenStudy (maddy1251):

\[-3x^{2}-6x-12\]

OpenStudy (maddy1251):

When I graphed this, my line was all the way down by (0,-12)

OpenStudy (maddy1251):

So, we will proceed with the next equation

OpenStudy (maddy1251):

\[ y = −3(x + 1)^{2} + 9\]

OpenStudy (maddy1251):

\[(x+1)^{2}=(x+1)(x+1)=x^{2}+1x+1x+1\]

OpenStudy (maddy1251):

\[x^{2}+2x+1\]

OpenStudy (maddy1251):

So now\[y=-3(x^{2}+2x+1)+9\]

OpenStudy (maddy1251):

Pretty sure that's it. Anyways, distribute -3

OpenStudy (maddy1251):

\[-3x^{2}-6x-3+9\]

OpenStudy (some.random.cool.kid):

yes thats it your clarified my answer so it was that :D. thanks.

OpenStudy (maddy1251):

\[y=-3x^{2}-6x+6\]

OpenStudy (maddy1251):

Would you look at that. A negative slope, the x is squared to give us the parabola, and we have the '+6' to verify our y-intercept

OpenStudy (maddy1251):

We can graph this now.

OpenStudy (some.random.cool.kid):

cool ok

OpenStudy (maddy1251):

Boom, there is our answer!

OpenStudy (maddy1251):

B is your answer.

OpenStudy (some.random.cool.kid):

yes it was right :~> :`> Ok next one way less tough and I believe that this is B also if were talking simpliest form. ofcourse I could be wrong. What is the equation, in vertex form, of the quadratic function that has a vertex at V(6, −7) and passes through point P(4, −9)? y equals start fraction one over two end fraction left parenthesis x plus six right parenthesis squared minus seven y = −2(x − 6)2 − 7 y equals negative start fraction one over two end fraction left parenthesis x minus six right parenthesis squared minus seven y equals start fraction one over two end fraction left parenthesis x plus six right parenthesis squared minus seven

OpenStudy (maddy1251):

Do you know what vertex form is?

OpenStudy (some.random.cool.kid):

wait dint read that yet.. its to messy i can separate this real quick.

OpenStudy (some.random.cool.kid):

y equals start fraction one over two end fraction left parenthesis x plus six right parenthesis squared minus seven y = −2(x − 6)2 − 7 y equals negative start fraction one over two end fraction left parenthesis x minus six right parenthesis squared minus seven y equals start fraction one over two end fraction left parenthesis x plus six right parenthesis squared minus seven there much better.

OpenStudy (some.random.cool.kid):

and no I am not sure im familiar with vertex yet. Well in math terms no the regular meaning yes.